OpenStudy (anonymous):

C=4d^-1/3b. Solve this formula for d. Any help is greatly appreciated :)

OpenStudy (anonymous):

\[\huge C=4d^{-\frac{1}{3}}b\]right?

OpenStudy (anonymous):

yep solvng for d

OpenStudy (anonymous):

move 4 and b to the other side. then log both sides

OpenStudy (anonymous):

I followed an example and did as they did and go tthe following answer : d=64b^3/c^3

OpenStudy (anonymous):

ya that's correct

OpenStudy (anonymous):

Theres was like this : D=1.1g^-1/4H. Solve for g D/1.1H = 1.1g^-1/4H/1.1H D^-4/(1.1H)^-4 = (g^-1/4)^-4 1.4641H^4/D^4=g

OpenStudy (anonymous):

ok Thank you ! I didnt see your reply earlier

OpenStudy (anonymous):

no problems

OpenStudy (anonymous):

I got c=1.89

OpenStudy (anonymous):

that's for solving for c

OpenStudy (anonymous):

how do you move the 4 and b on the other side, do you divide or subtract

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

kk